OpenStudy (anonymous):

Please solve (r-3)(r+6)=0

OpenStudy (anonymous):

r - 3 =0 r - 3 + 3 = 0 + 3 r = 3 r + 6 = 0 r + 6 - 6 = 0 - 6 r = -6 r = 3, -6

OpenStudy (anonymous):

set each parenthesis against 0 r-3=0 r=3 r+6=0 r=(-6)